OpenStudy (anonymous):

In the below system, solve for y in the first equation. x – 3y = –6 2x – 7y = 10 x + 2 x + 6 –x + 2 x + 2 i think its A. but im not sure...

OpenStudy (anonymous):

these r the real answer choices my bad a. 1/3x+2 b. -1/3x+6 c. -x+2 d. -1/3x+2

OpenStudy (ivettef365):

you have equation x -3y = -6 -3y = -x -6 (I transfer the x to this side changing the sign) --- --- -- -3 -3 -3 (divide everything by -3) then you have y = 1/3x +2

OpenStudy (anonymous):

so i was right?

OpenStudy (ivettef365):

correct